Why is the answer $12?

Why is the answer D?
Answer #1

1. Option D. $12

Explanation: Consumer surplus is the difference between the maximum willingness to pay and the price. Here, only 4 units would be sold as the willingness to pay for the 5th unit is lower than the price.

Total consumer surplus = sum of the consumer surplus of individual products = (14 - 8) + (12-8) + (10-8) + (8-8) = 6 + 4 + 2 + 0 = 12.

2. Option D

Explanation: When the price floor is P1, it is below the equilibrium price. A price floor below the equilibrium price is non-binding and it does not affect price or quantity i..e there is no deadweight loss.